OpenStudy (anonymous):

Find the exact value by using a sum or difference identity. cos 19π/12

OpenStudy (anonymous):

god am i wrong today. lets try again

OpenStudy (anonymous):

\[\frac{19}{12}=\frac{5}{4}+\frac{1}{3}\]

OpenStudy (anonymous):

that is what i want. so you have \[\cos(\frac{19\pi}{12})=\cos(\frac{5\pi}{4}+\frac{\pi}{3})\]

OpenStudy (anonymous):

now use addition angle formula for cosine. are you good from there or do you need it worked out?